# femto-PIXAR: a neural network method for reconstruction of femtosecond X-ray free electron laser pulse energy

### v1

@@ Abstract
X-ray Free Electron Lasers (X-FELs) operate in a wide range of lasing configurations for a broad variety of scientific applications at ultrafast time-scales such as structural biology, materials science, and atomic and molecular physics. Shot-by-shot characterization of the X-FEL pulses is crucial for analysis of many experiments as well as tuning the X-FEL performance, but an outstanding challenge exist for monochromatic multi-bunch studies, such as those needed for the study of fluctuations in quantum materials. In particular, there is no current method for reliably resolving the pulse intensity profile at sub-picosecond time separation between pulses. Here we show that a physics-based U-net model can reconstruct the individual pulse power profiles on the sub-picoseond pulse separation time. Using experimental data from weak X-FEL pulse pairs, we demonstrate we can learn the pulse characteristics on a shot-by-shot basis when conventional methods fail.
@@ I Introduction
We address this need by developing a novel self-supervised machine learning (ML) framework called femto-PIXAR: femtosecond Photon Inference of X-ray pulses using AI References, which provides sufficient sensitivity to resolve microjoule-scale pulses with femtosecond pulse separation. We show our approach accurately predicts pulse energies for an experimental dataset, verified against total pulse energy monitors in real X-FEL data, and confirm the predictions match the expected pulse delays. This demonstration will make fluctuation studies possible in the sub-picosecond regime, but will also provide a tool for a wide array of X-FEL experiment in many other fields as well.
@@ II Setup
The experimental setup was designed to produce two coherent X-ray pulses starting from a single, long electron bunch, itself generated from two overlapping injector lasers. The two X-ray pulses were created using a double-slotted V-shaped foil (for time domain control) [18] [19] and hard X-ray self-seeding [20] (for coherence and energy selection). To our knowledge this configuration, described in Fig. 1 a), has never been used before. The experiment was conducted at a photon energy of 8.35 keV (Ni -edge), with slotted-foil pulse separations of (single pulse), , and . Different stages of the electron bunch shaping sequence are labeled , , , , representing the same electron distribution at different instances as it proceeds in time. A 3 m thick aluminum foil with two slots of variable separation from 1 mm to 5 mm is placed in the middle of the dispersive section of the bunch compressor BC2 [18]. At the locations where the electron bunch passes through the foil the electron emittance is ‘spoiled’ (i.e. increases) due to coulomb scattering. The emittance growth suppresses FEL amplification in the undulator, so that only the electrons that pass through the slits of the V-shaped foil produce X-ray pulses.
Downstream of the bunch compressor the relative distance between the unspoiled electron distributions is now in the direction of travel, resulting in an arrival time difference , that can be changed by shifting the slotted foil. passes through the undulator magnets resulting in X-ray lasing. Between the first and second undulator segment a hard X-ray self-seeding scheme is applied (not shown here). Downstream of the undulators, the X-band Transverse deflecting mode CAVity (XTCAV) is employed. [21]. The deflected electron bunch is streaked across the target screen to image the longitudinal distribution, while the X-ray gas-monitor detector (XGMD) [22] measures the total photon pulse energy. The X-rays continue down the beamline to the experimental endstations. This entire process repeats at a 120 Hz repetition rate.
@@ III Data and Methodology
The XTCAV images the longitudinal (time-energy) distribution of the post-lasing electron bunch. By comparison with a lasing-off reference, it is possible to infer the temporal profile of the photon pulse. However, because the lasing-off reference cannot be measured simultaneously, the reference must be inferred from the lasing-on measurement. The “classical” approach has been to record a set of images while suppressing lasing, cluster these images by shape, and then select the most similar cluster for each lasing-on image by comparing the projected electron current [21, 23, 24]. While the classical approach has been demonstrated for various two-pulse modes [19, 25, 14], it struggles for small pulse energies, when shot-to-shot differences between the true lasing-off electron bunch and the selected reference cluster exceed the lasing-induced changes. As a result, when combining seeding with two-pulse modes, the classical method breaks down.
@@
A deep learning method proposed by Ren et al. [26] sidesteps the need for matched lasing-off references by training a neural network (NN) directly to predict power from a lasing-on image; however, the NN relies on accurate simulations for training, and the lack of transverse phase space diagnostics makes this particularly challenging for the complex distributions in the conditions implementing the slotted foil. On the other hand, for the slotted foil setup, only a narrow segment of the beam lases. We can exploit this feature to approximate a reference in the short lasing region by using information in the nearby non-lasing regions. For example, Zeng et al. [27] suggested to use a polynomial regression across the lasing region as a reference. However, the polynomials cannot model the complex phase space of this setup and fail to reproduce the weak lasing pulses (see Fig. 3).
@@
Here, we instead propose an approach using a U-net [28] to generate the corresponding lasing-off reference image. We train the NN using self-supervision, eliminating the need for simulations. Specifically, we mask regions of lasing-off pulses, and train the NN to regenerate those regions. Once trained, the network receives a lasing-on image with masked lasing regions and reconstructs the best lasing-off reference. We then calculate the center of mass (COM) to retrieve the X-ray pulse distribution [24]. The network can also detect the lasing region by identifying areas with the largest reconstruction discrepancy. Our method provides high-fidelity lasing-off references without relying on the accuracy of high-performance computing simulations while retaining interpretability through direct observation of the generated reference.
@@
Figure 1 b) and c) show the core steps. The input of the network are cropped, background subtracted and normalized lasing off images, where all values within two random mask-regions are set to zero (black bars in Fig. 1 b) and c)). The width of an individual mask was chosen to be 16 pixels and the distance range for the two masks was set from 0 pixel to 40 pixel, corresponding to approx. 0 to 38 fs. We then train a U-net to reconstruct the electron phase space of the lasing-off data (Fig. 1 b) ). As the center of mass is the critical value for evaluation, we use a loss that combines the Mean Squared Error of pixel values with the absolute error of the COM. After training, we can mask the lasing regions in a lasing-on data set. As the network is trained only on lasing-off regions of phase space, it will reconstruct the best matching lasing-off electron phase space exactly for a given image (Fig. 1 c). This lasing-off image will also exactly match the original x and y position, so no shifting in the time axis or normalization with the XGMD (as in the classical approach) is necessary.
@@
A crucial step is this process is in determining correct mask positions in lasing-on data. The separation of the two pulses is approximately fixed by the slit separation, but due to shot-to-shot differences in electron bunch energy and compression of the electron bunch, the position of the features relative to the center of mass of the electron bunch is not constant. Our method scans the image to find two locations where input and reconstruction differ most, centering the masks on the maximum signal. We then use the methods from Behrens et. al, marked as in Fig. 1[24] to obtain the power profile using the spectral COM from both images.
@@ IV Results
Fig. 2 shows example reconstructions for the three different configurations and one lasing-off example, all randomly selected from the top 1 of data with the highest XGMD values (ensuring a clearly visible signal). None of these images were part of the network training process. To demonstrate the efficacy of the proposed method, we benchmark the predicted power and delay against existing diagnostics. First, we compare the integral of the predicted profile with the total energy of both pulses measured by the XGMD, which cannot resolve pulses on the fs scale. In Fig. 3 one can see the correlation of the integral of the power profile with the XGMD for the three different slotted foil configurations. We note that while there is a strong linear correlation between the predicted power and XGMD monitor, there is a disagreement in proportionality. Further experiments are necessary to determine if it is due to either the XGMD or XTCAV calibration, or bias from the method. However, as the typical user experiment usually relies primarily on the relative power between the pulse pairs, this discrepancy does not affect the experimental analysis.
We compare our predictions to two existing state-of-the-art XTCAV analysis methods. In order to evaluate the correlation with the XGMD, in the classical approach (see the previous section), the energy jitter (shift in the y-axis) cannot be compensated for by normalization with the XGMD but must instead be corrected using a scheme that matches the head and tail of the COM profiles for the lasing-on and lasing-off images. We also compare to the polynomial regression method of Zeng et al. [27]. Some example predictions of this method can be found in the supplemental material. Figure 3 compares the XGMD correlation for all three delays, neither of which results in the expected correlation.
Fig. 4 shows the temporal and energy characteristics of the X-ray laser pulses as retrieved by the U-net NN reconstruction scheme. The time assignments of each shot are obtained by fitting the X-ray power trace with two Gaussian functions. For these plots we take only the data where both gaussian curves indicate a signal larger than 15 J. We find that the predicted time difference between both pulses closely matches the delay given by the slotted foil condition. The temporal separation jitter aligns well with the 9 jitter observed by Ding et al. [19]. As an additional benchmark we checked the predicted pulse energies for lasing-off data with randomly placed masks and find that 80 of the data has a deviation from zero of less than 15 J, and 99.8 less than 25 J. We also find that the pulse energy difference between the two pulses is centered well around zero for lasing-off-data, indicating that our approach does not introduce a bias in the ratio between the first (left) and the second (right) pulse (see supplemental material).
@@ V Discussions and Summary
In conclusion, femto-PIXAR is a novel X-FEL diagnostic which we demonstrate by reconstructing the X-ray power profiles from weak X-FEL pulses in a non-standard two-pulse, self-seeding configuration. For this configuration, the classical method of XTCAV analysis, which uses a matching algorithm to subtract similar lasing-off shots for the energy loss calculation, fails to reproduce meaningful correlations with the X-ray gas detector. By contrast, our method is able to distinguish the relative amplitude of each X-ray pulse on a shot-to-shot basis which is necessary for X-ray photon fluctuation spectroscopy experiments and other future experiments at X-FEL facilities. Our approach has various benefits, avoiding the need for energy calibration with an external reference and scaling well for high-throughput experiments enabled by next generation X-FEL facilities such as the LCLS-II [29]. As opposed to previous deep-learning work, our approach uses self-supervision with a physics-motivated loss function, avoiding the need for high-fidelity simulations. Our method also reconstructs the full 2D phase space, providing interpretability of the network and giving crucial information about when the network might not be applicable for a specific dataset. Future work will consider extensions to more general XFEL configurations beyond the slotted foil for generation of sub-picosecond-spaced pulses. For example, using diagnostics beyond the XTCAV could enable analysis of configurations with lasing across the full electron bunch [30]. The analysis chain after reconstructing the lasing-off reference can also make use of machine learning, e.g. incorporating spectral measurements to refine the x-ray pulse profile [31].

### v2

# femto-PIXAR: a self-supervised neural network method for reconstructing femtosecond X-ray free electron laser pulses

@@ Abstract
X-ray Free Electron Lasers (X-FELs) operate in a wide range of lasing configurations for a broad variety of scientific applications at ultrafast time-scales such as structural biology, materials science, and atomic and molecular physics. Shot-by-shot characterization of the X-FEL pulses is crucial for analysis of many experiments as well as tuning the X-FEL performance. However, for the weak pulses found in advanced configurations, e.g. those needed for coherent, two-pulse studies of quantum materials, there is no current method for reliably resolving pulse profiles. Here we show that a physics-based U-net model can reconstruct the individual pulse power profiles for sub-picosecond pulse separation without the need for simulations. Using experimental data from weak X-FEL pulse pairs, we demonstrate we can learn the pulse characteristics on a shot-by-shot basis when conventional methods fail.
@@ I Introduction
We address this need by developing a novel self-supervised machine learning (ML) framework called femto-PIXAR: femtosecond Power Inference of X-ray pulses using AI References, which provides sufficient sensitivity to resolve microjoule-scale pulses with femtosecond pulse separation. We show our approach accurately predicts pulse energies for an experimental dataset, verified against total pulse energy monitors in real X-FEL data, and confirm the predictions match the expected pulse delays. This demonstration will make fluctuation studies possible in the sub-picosecond regime, and will also provide a tool for a wide array of X-FEL experiments in many other fields as well.
@@ II Setup
The experimental setup was designed to produce two coherent X-ray pulses starting from a single, long electron bunch, itself generated from two overlapping injector lasers. The two X-ray pulses were created using a double-slotted V-shaped foil (for time domain control) [18] [19] and hard X-ray self-seeding [20] (for coherence and energy selection). To our knowledge this configuration, described in Fig. 1 a), has never been used before. The experiment was conducted at a photon energy of 8.35 keV (Ni -edge), with slotted-foil pulse separations of (single pulse), , and . Different stages of the electron bunch shaping sequence are labeled , , , , representing the same electron distribution at different instances as it proceeds in time. A 3 m thick aluminum foil with two slots of variable separation from 1 mm to 5 mm was placed in the middle of the dispersive section of the bunch compressor BC2 [18]. At the locations where the electron bunch passes through the foil the electron emittance was ‘spoiled’ (i.e. increased) due to coulomb scattering. The emittance growth suppresses FEL amplification in the undulator, so that only the electrons that pass through the slits of the V-shaped foil produce X-ray pulses.
Downstream of the bunch compressor the relative distance between the unspoiled electron distributions is in the direction of travel, resulting in an arrival time difference , that can be changed by shifting the slotted foil. passes through the undulator magnets resulting in X-ray lasing. Between the first and second undulator segment a hard X-ray self-seeding scheme is applied (not shown here). Downstream of the undulators, the X-band Transverse deflecting mode CAVity (XTCAV) streaks the electron beam on a target to image the final longitudinal phase space [21]. The X-ray laser pulses are unaffected by the XTCAV deflector and continue downstream to the X-ray gas energy monitor (GEM), where the total X-ray pulse energy is transmissively monitored on a shot-to-shot basis using a photomultiplier tube (PMT) to measure the fluorescence induced in a small volume of N2 gas [22]. After the GEM, the X-rays continue down the beamline to the experimental endstations. This entire process repeats at a 120 Hz repetition rate.
@@ III Data and Methodology
The XTCAV images the longitudinal (time-energy) distribution of the post-lasing electron bunch. By comparison with a lasing-off reference, it is possible to infer the temporal profile of the photon pulse. However, because the lasing-on and lasing-off references cannot be measured simultaneously, this poses a fundamental challenge in determining the change the electron phase space due to the undulators. The “classical” approach has been to record a set of images while suppressing lasing, cluster these images by shape, and then select the most similar cluster for each lasing-on image by comparing the projected electron current [21, 23, 24]. While the classical approach has been demonstrated for various two-pulse modes [19, 25, 14], it struggles for small pulse energies, when shot-to-shot differences between the true lasing-off electron bunch and the selected reference cluster exceed the lasing-induced changes. As a result, when combining seeding with two-pulse modes, the classical method breaks down.
@@
A deep learning method proposed by Ren et al. [26] sidesteps the need for matched lasing-off references by training a neural network (NN) directly to predict power from a lasing-on image; however, the NN in [26] relies on simulated labels for training, and it is particularly challenging to simulate the complex electron beam distribution produced by the slotted foil. On the other hand, for the slotted foil setup, only a narrow segment of the beam lases. We can exploit this feature to approximate a reference in the short lasing region by using information in the nearby non-lasing regions. For example, Zeng et al. [27] suggested to use a polynomial regression across the lasing region as a reference. However, the polynomials cannot model the complex phase space of this setup and fail to reproduce the weak lasing pulses (see Fig. 3).
@@
Here, we instead propose an approach using a U-net [28] to generate the corresponding lasing-off reference image. We trained the NN using self-supervision, eliminating the need for simulations. Specifically, we masked regions of lasing-off pulses, and trained the NN to regenerate those regions. Once trained, the network receives a lasing-on image with masked lasing regions and reconstructs the best lasing-off reference. We then calculated the center of mass (COM) to retrieve the X-ray pulse distribution [24]. The network can also detect the lasing region by identifying areas with the largest reconstruction discrepancy. Our method provides high-fidelity lasing-off references without relying on the accuracy of computationally-expensive simulations while retaining interpretability through direct observation of the generated reference.
@@
Fig. 1 b) and c) shows the core steps. The inputs to the network were cropped, background subtracted and normalized lasing off images, where all values within two random mask-regions were set to zero (black bars in Fig. 1 b) and c)). The width of an individual mask was chosen to be 16 pixels and the distance range for the two masks was set from 0 to 40 pixels, corresponding to approx. 0 to 38 fs. We then trained a U-net to reconstruct the electron phase space of the lasing-off data (Fig. 1 b) ). As the center of mass is the critical value for evaluation, we used a loss that combines the Mean Squared Error of pixel values with the absolute error of the COM. After training, we masked the lasing regions in a lasing-on data set. As the network is trained only on the lasing-off data, it reconstructs the best matching lasing-off electron phase space exactly for a given image (Fig. 1 c). This lasing-off image will also exactly match the original x and y position, so no shifting in the time axis or normalization with the GEM (as in the classical approach) is necessary.
@@
A crucial step of this process was to determine the correct mask positions in the lasing data. The separation of the two pulses was approximately fixed by the slit separation, but due to shot-to-shot differences in electron bunch energy and compression of the electron bunch, the position of the features relative to the center of mass of the electron bunch was not constant. Our method scans the image to find two locations where input and reconstruction differ most, centering the masks on the maximum signal. We then use the methods from Behrens et. al [24], marked as in Fig. 1 to obtain the power profile using the spectral COM from both images.
@@ IV Results
Fig. 2 shows example reconstructions for the three different configurations and one lasing-off example, all randomly selected from the top 1 of data with the highest GEM values (ensuring a clearly visible signal). None of these images were part of the network training process. To demonstrate the efficacy of the proposed method, we benchmark the predicted power and delay against existing diagnostics. First, we compare the integral of the predicted profile with the total energy of both pulses measured by the GEM, which cannot resolve pulses on the fs scale. In Fig. 3 one can see the correlation of the integral of the power profile with the GEM for the three different slotted foil configurations. We note that while there is a strong linear correlation between the predicted power and GEM monitor, there is a disagreement in proportionality. Further experiments are necessary to determine if it is due to either the GEM or XTCAV calibration, or bias from the method. However, as the typical user experiment usually relies primarily on the relative power between the pulse pairs, this discrepancy does not affect the experimental analysis.
We compare our predictions to two existing state-of-the-art XTCAV analysis methods. In order to evaluate the correlation with the GEM, in the classical approach (see the previous section), the energy jitter (shift in the y-axis) cannot be compensated for by normalization with the GEM but must instead be corrected using a scheme that matches the head and tail of the COM profiles for the lasing-on and lasing-off images. We also compare our results with the polynomial regression method of Zeng et al. [27]. Some example predictions of this method can be found in the supplemental material. Fig. 3 compares the GEM correlation for all three delays, neither of which results in the expected correlation.
Fig. 4 shows the temporal and energy characteristics of the X-ray laser pulses as retrieved by the U-net reconstruction scheme. The time assignments of each shot were obtained by fitting the X-ray power trace with two Gaussian functions. For these plots we took only the data where both Gaussian curve fits indicated a signal larger than 15 J. We find that the predicted time difference between both pulses closely matches the delay given by the slotted foil condition. The temporal separation jitter aligns well with the 9 jitter observed by Ding et al. [19]. As an additional benchmark we checked the predicted pulse energies for lasing-off data with randomly placed masks and find that 80 of the data has a deviation from zero of less than 15 J, and 99.8 less than 25 J.
@@ V Discussions and Summary
In conclusion, femto-PIXAR is a novel X-FEL diagnostic which we demonstrate by reconstructing the X-ray power profiles from weak X-FEL pulses in a non-standard two-pulse, self-seeding configuration. For this configuration, the classical method of XTCAV analysis, which uses a matching algorithm to subtract similar lasing-off shots for the energy loss calculation, fails to reproduce meaningful correlations with the X-ray gas detector. By contrast, our method is able to distinguish the relative amplitude of each X-ray pulse on a shot-to-shot basis which is necessary for X-ray photon fluctuation spectroscopy experiments and other future experiments at X-FEL facilities. Our approach has various benefits, avoiding the need for energy calibration with an external reference. A neural network based scheme also scales well for the large datasets expected from high-throughput experiments enabled by next generation X-FEL facilities such as the LCLS-II [29]. As opposed to previous deep-learning work, our approach uses self-supervision with a physics-motivated loss function, avoiding the need for high-fidelity simulations. Our method also reconstructs the full 2D phase space, providing interpretability of the network and giving crucial information about when the network might not be applicable for a specific dataset. Future work will consider extensions to more general X-FEL configurations beyond the slotted foil. For example, using diagnostics beyond the XTCAV could enable analysis of configurations with lasing across the full electron bunch [30], e.g. to provide the femtosecond resolution needed for temporal ghost imaging [31]. The analysis chain after reconstructing the lasing-off reference can also make use of machine learning, e.g. incorporating spectral measurements to refine the X-ray pulse profile [32].